OpenStudy (taintedflame):

A parachutist's rate during a free fall reaches 65 meters per second. What is this rate in feet per second? At this rate, how many feet will the parachutist fall during 10 seconds of free fall? In your computations, assume that 1 meter is equal to 3.3 feet. Do not round Rate:___ ft/s Distance fallen in 10 seconds:___ ft

OpenStudy (taintedflame):

last wuestion on my homework and im stumped

OpenStudy (sshayer):

this distance is covered in 1 s multiply by 10 to find distance covered in 10 seconds.

OpenStudy (sshayer):

1m=3.3 ft 65 m=65*3.3=?

OpenStudy (niksbhalla14mar):

As speed is 65 m / sec so Speed would be (65x 3.3) ft/sec = 195 ft/sec In free falling case. gravity is 9.8 m/sec2 which means that speed is changing by 9.8m/sec in one second. And Distance dropped is calculated by h=1/2 gt^2. So distance in 10 sec would be 490 m.